OpenStudy (anonymous):

Given f(x) = 8x + 13 and g(x) = x2 - 5x, find the function (f - g)(x). Use the ^ key for the exponent. Enter the answer as the following example: -5x^2+4x-8

OpenStudy (gorv):

(f-g)(x)=f(x)-g(x) =8x+13-x^2+5x =-x^2+8x+5x+13 rest u can do ithink